# Keep the Cost Down: A Review on Methods to Optimize LLM’s KV-Cache Consumption

### v2

@@ 5 Post-Training Optimizations
5.1 Eviction
While the memory for KV-Cache is limited, LLMs have to process long context, resulting in memory overhead. One solution is to keep a fixed-size of KV-Cache, storing the critical token Key-Value (KV) pairs. Eviction methods are about the policies to discard unnecessary tokens. Two lines of approaches exist: static policies, which are designed manually, and dynamic policies, which utilizes attention scores or other information to identify important tokens.
Static policies A straightforward approach for maintaining a fixed-size KV-Cache is to retain recent tokens, known as window attention (Beltagy et al., 2020). However, the model will collapse when sequence length exceeds cache size (Xiao et al., 2023). Recently, Xiao et al. (2023) and Han et al. (2024) suggest that initial tokens almost consistently receive high attention weights across layers and heads. This indicates that removing them significantly impacts the softmax function calculation. As a result, keeping KV-Cache of both initial and recent tokens can maintain model performance for long contexts.
@@ 5.1 Eviction
This provides us with the possibilities to estimate future significance of a token based on history information of attention weights. Token Omission Via Attention (TOVA) (Oren et al., 2024) proposes a simple greedy method to discard useless tokens. While keeping a fixed-length of KV-Cahce, TOVA evicts at each decoding step the tokens with minimal attention weights layer-wise when using the last token to queries all past keys and values. H2 Eviction Algorithm (Zhang et al., 2023) uses accumulative normalized attention scores to decide which token to stay and at the same time keeps the recent tokens since they may show strong correlations with current tokens. Liu et al. (2023) uses a counter to keep track of the number of times when a certain token is regarded as insignificant using attention weights as a criterion. This algorithm also keeps the recent tokens since a lack of information about their significance. However, while discarding unnecessary tokens seems to have minor influence on the original inference process, Adnan et al. (2024) observes that when evicting more tokens during the inference process, the distribution of normalized attention scores becomes uneven among the remaining ones. Therefore, Keyformer(Adnan et al., 2024) introduces an additional distribution to smooth the uneven distribution and approximate the original one when using full KV-Cache. It also introduces a temperature factor during decoding process. With the advances in decoding process, the temperature becomes greater, resulting in more randomness and smooth distribution. FastGen (Ge et al., 2023) uses a hybrid strategy to optimize token selection for discarding. In prompt encoding phase, it selects the best policy for each head in KV-Cache, and then uses these policies to decide which token to discard in decoding process. The policies include: a) keeping special tokens, b) keeping punctuation tokens, c) keeping recent tokens and d) keeping tokens with attention-weight-based policies. SparQ Attention (Ribar et al., 2023) tries a different approach. Instead of reducing memory capacity, it aims at reducing the amount of data transferred. It first uses the norm of $q_{i}^{(l)}$ to decide which $k_{i}^{(l)}$ needs to be fetched, and approximate the attention scores based on this subset of queries and keys. Then it selects the top-$K$ attention weights and fetch their corresponding Key-Value pair to calculate the output. To compensate the values that are considered insignificant, it keeps a running mean of value vectors and interpolates between this running mean and the output of attention so as to approximate the original distribution.
@@ 5.2 Quantization
Another commonly utilized method for compressing Key-Value (KV) cache is through quantization. This approach effectively compresses data by mapping tensor values, originally in full precision, to discrete levels and storing them at a reduced precision. There are two primary categories of KV-Cache quantization: Full quantization and KV-Cache-only quantization. Full quantization involves compressing both the model weights and the KV-Cache, thereby reducing the memory footprint of the entire model. On the other hand, KV-Cache-only quantization specifically targets the KV-Cache activations, selectively compressing them to conserve memory while leaving the model weights in their original state. This tailored approach can offer a balance between model efficiency and performance.
Figure 3: The comparison between different KV-Cache policies
KV-Cache only quantization
@@
Dong et al. (2024a) introduces LESS (Low-rank Embedding Sidekick with Sparse policy), which integrates a constant-sized cache with eviction-based cache methods (Section 5). Inspired by recurrent networks, LESS utilizes a constant-sized low-rank cache by replacing the softmax with a separable similarity metric calculated by learnable row-wise functions. By doing so, LESS obtains the ability to accumulates history token information before they are discarded from KV-Cache, allowing for continued access to previous information.
Yang et al. (2024) proposd Mixed-precision KV-Cache (MiKV), a reliable cache compression method that retains evicted KV pairs in reduced precision to preserve information and maintains important KV pairs in higher precision to ensure generation quality. MiKV employed Dynamic Outlier Awareness to dynamically balance the outliers manifested in the query and keys to reduce quantization error. Dynamic Outlier Awareness multiplies and divides a channel balancer to the keys and queries to mitigate the impact of outliers.
@@
Yue et al. (2024) took FlexGen a step further and proposed WKVQuant. For weight quantization, WKVQuant ultize the OmniQuant (Shao et al., 2023). For KV-Cache quantization, WKVQuant employs following three strategies: Past Only Quantization, Two-dimensional Quantization and Cross-block Reconstruction Regularization. Different from KVQuant (Hooper et al., 2024) which focus on calibrating for per-channel(token) thresholds to ignore outlier, WKVQuant focuses on aligning and smoothing each channel and token.
@@ 6 Evaluation
In this section, we first introduce the datasets commonly used in KV-Cache optimization, followed by a presentation of frequently used evaluation metrics.
@@ 6.1 Datasets
Longbench
LongBench (Bai et al., 2023) is the first bilingual (English and Chinese) multitask benchmark for long context understanding. It consists of 21 datasets across 6 task categories: single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks, and code completion. The average length of instances is 6,711 words for English and 13,386 characters for Chinese.
Passkey retrieval
Mohtashami & Jaggi (2023) proposed the passkey retrieval task. In this task, the models are required to retrieve a random passkey hidden in a long document. An example for passkey retrieval task can be found in Appendix B Figure 4
Needle in a Haystack
Needle in a Haystack task is first proposed by (Kuratov et al., 2024). They proposed BABILong, a new benchmark designed to assess model capabilities in extracting and processing distributed facts within extensive texts. BABILong hides algorithmically generated question answering and reasoning problems inside a corpus of book texts. BABILong consists of 20 tasks designed for evaluation of basic aspects of reasoning. An example for Needle in a Haystack task can be found in Appendix B Figure 5.
@@
In addition to datasets specifically designed for long texts, the length of traditional shorter test sets can be extended through a few-shot (Brown et al., 2020) format or by simulating multi-turn dialogues, in order to test the model’s capabilities with long texts. Furthermore, for some inference-type tests, the Chain-of-Thought (CoT) strategy proposed in Wei et al. (2022) can be adopted to further increase the length of few-shot texts. However, in long-text tests based on few-shot, texts at a greater distance only serve as a reference and do not contain information that the model needs to extract or conditions for inference, so the reference value of the results is limited.
@@ 6.2 Evaluation Metric
For KV-Cache, the most intuitive optimization indicator would be the memory space occupied by each token. The LLaMA2-7B model, as a typical example, theoretically occupies $0.5$MB of memory for each KV-Cache entry, while the structurally similar but MQA-using Mistral-7B only occupies $0.125$MB, which is a quarter of the former. Please note that when measuring this indicator rigorously, the fragment space generated by the token should also be taken into account, that is, it is best to measure according to the actual memory occupancy, rather than a value calculated through a string of parameters.
@@
Throughput and latency are important indicators for measuring the time efficiency of a model. Throughput, usually measured in tokens per second (token/s), represents how many new tokens the model can generate per second. The higher the throughput, the better the model performance. There are more possibilities for the definition of latency. For the pre-fill phase, latency can be represented as the time required to process all tokens, usually in milliseconds, and is generally similar for sequences of the same length. In the Decoding phase, latency is usually considered to be the time required to generate each new token, which is the reciprocal of the throughput, typically in milliseconds.
@@ 7 Key Takeaways
This review, following the footsteps of prior work, takes a closer look into KV-Cache optimization, uncovering its complexities and proposing strategies for its optimization. We hope to propose some insights and suggested directions for future research that are not just reflections of current trends but also an invitation to explore the uncharted territory of LLMs.
Principles of KV-Cache Optimization: At the core of optimizing KV-Cache lies the principle of reducing memory consumption. This can be achieved by further compressing ’K’ (Keys) or ’V’ (Values) in the KV pairs. Techniques to compress these components directly impact the efficiency of the models, especially in terms of memory usage and processing speed.
Trade-offs in Deletion vs. Compression: Whether to delete less important KV pairs to save memory or to compress the KV-Cache without deletion remains an open question. While deletion might offer immediate memory relief, it could potentially compromise the model’s performance. In contrast, better compression techniques strive to retain information integrity while reducing memory footprint.
Extremes in KV-Cache Management: A more radical approach could involve storing the KV-Cache externally, possibly on a different storage medium. This method would transform KV-Cache management into a retrieval challenge, where the relevant KV pairs are fetched and reintegrated into the model as needed. While this could reduce memory usage on primary devices, it would introduce complexities in retrieval and integration processes.
Future Directions in Storage and Retrieval Technologies: These discussions point towards an evolving future where storage and retrieval technologies might become as crucial as the computational models themselves. Innovations in how KV-Cache is managed, stored, and accessed could open up new avenues for making LLMs more efficient and versatile.
@@ 8 Conclusion
In this review, we highlighted the significance and the multifaceted nature of optimizing KV-Cache in Large Language Models (LLMs). This review traversed through a variety of methods, from the training and deployment stages to post-training optimizations. Each method, whether it be the architectural changes in the pre-training phase, framework optimizations during deployment, or dynamic strategies like eviction and quantization in the post-training phase, offers unique insights into mitigating the challenges posed by the KV-Cache’s memory-intensive nature. Moreover, the exploration of different evaluation metrics for long-text performance underpins the importance of a balanced approach that considers both efficiency and model capabilities. As the field continues to evolve, it is clear that the optimization of KV-Cache will remain a critical area of focus, offering a pathway towards more efficient and environmentally responsible use of LLMs. We hope this review can be a roadmap to guide learners venturing through this dynamic and rapidly progressing field.
@@ Appendix A Popular Models With GQA
  • $S_{HK}$ and $S_{KV}$: Embedding size (in bytes, bfloat16 as data type) and corresponding KV-Cache increment per token.
@@
  • $\mathcal{R}$: Represents the ratio of the KV-Cache size generated for each token to the size of the embedding vector. A higher ratio indicates that the information for this token has been expanded by a greater factor. Notably, when using the GQA method, this value is significantly lower compared to non-GQA methods, implying that the information produced is relatively compressed.
@@ Appendix C Other Methods
In addition to the above methods, there are also some methods that modify the Attention mechanism to completely eliminate KV-Cache. Furthermore, by directly compressing the prompt or the embedding, the length of KV-Cache can also be reduced by decreasing the number of tokens.
@@ C.1 Linear-Transformers
Katharopoulos et al. (2020); Wang et al. (2020); Verma (2021) proposed a linear attention mechanism. By removing the softmax function and slightly adjusting $W_{\{q,k,v\}}$, the computational complexity of the Attention part in the Transformer is transformed from quadratic to linear, thereby achieving an inference speed consistent with Recurrent Neural Networks (RNNs), and totally eliminate the KV-Cache.
Zhang et al. (2024); Arora et al. (2024) are some improvements to linear attention. These include the use of naive Attention mechanisms at close range, and the optimization of the performance of linear Transformers through Spiky and Monotonic Weights.
@@ C.2 Prompt and Embedding Engineering
LLMLingua (Jiang et al., 2023b) uses a budget controller, an iterative token-level compression algorithm and a small model to estimate which token is important. LongLLMLingua (Jiang et al., 2023c) further improves LLMLingua in long context scenarios.

### v3

@@ 5 Post-Training Optimizations
5.1 Eviction and Merging
Eviction methods are about the policies to discard unnecessary tokens. Two lines of approaches exist: static policies, which are designed before inference and remains consistent across every inference request, and dynamic policies, which utilize the information generated while inference to identify important tokens.
Static policies A straightforward approach for maintaining a fixed-size KV Cache is to retain recent tokens, known as slide window attention (Beltagy et al., 2020). However, the model will collapse when sequence length exceeds cache size (Xiao et al., 2023). Recently, Xiao et al. (2023) and Han et al. (2024) proposed that initial tokens almost consistently receive high attention weights across layers and heads. As a result, keeping KV Cache of both initial and recent tokens can maintain model performance for long contexts.
@@ 5.1 Eviction and Merging
This provides us with the possibilities to estimate future significance of a token based on history information of attention weights. Token Omission Via Attention (TOVA) (Oren et al., 2024) proposes a simple greedy method to discard useless tokens. While keeping a fixed-length of KV-Cahce, TOVA evicts at each decoding step the tokens with minimal attention weights layer-wise.
H2 Eviction Algorithm (Zhang et al., 2023) uses accumulative normalized attention scores to decide which token to stay and at the same time keeps the recent tokens since they may show strong correlations with current tokens.
PyramidInfer(Yang et al., 2024a) uses a layer-wise approach with recent tokens occupying more weights. Apart from this, Yang et al. (2024a) believes that deeper layer contains more context redundancy, so it defines a decay ratio to shorten the length of KV Cache in deeper layers and thus forming a ”pyramid”.
However, while discarding unnecessary tokens seems to have minor influence on the original inference process, Adnan et al. (2024) observes that when evicting more tokens during the inference process, the distribution of normalized attention scores becomes uneven among the remaining ones. Therefore, Keyformer(Adnan et al., 2024) introduces additional methods to smooth the uneven distribution and approximate the original one when using full KV Cache.
FastGen (Ge et al., 2023) uses a hybrid strategy to optimize token selection for discarding. In prompt encoding phase, it selects the best policy for each head in KV Cache, and then uses these policies to decide which token to discard in decoding process. The policies include: a) keeping special tokens, b) keeping punctuation tokens, c) keeping recent tokens and d) keeping tokens with attention-weight-based policies.
SparQ Attention (Ribar et al., 2023) tries a different approach. Instead of reducing memory capacity, it aims at reducing the amount of data transferred. It first uses the norm of $q_{i}^{(l)}$ to decide which $k_{i}^{(l)}$ needs to be fetched, and approximate the attention scores based on this subset of queries and keys. Then it selects the top-$K$ attention weights and fetch their corresponding Key-Value pair to calculate the output. To compensate the values that are considered insignificant, it keeps a running mean of value vectors and interpolates between this running mean and the output of attention so as to approximate the original distribution.
Methods based on attention score also includes Corallo & Papotti (2024); rae Jo & Shin (2024); Xu et al. (2024); Fu et al. (2024).
While attention scores provide ample information, they are incompatible with certain optimization methods, such as Flash Attention (Shah et al., 2024). Therefore, investigating how to evict tokens without relying on attention scores is a novel research topic. Several approaches address this challenge. Tang et al. (2024) leverages the prior-known properties of attention heads to determine whether to evict. Yao et al. (2024) uses token entropy to decide whether to remove tokens. Devoto et al. (2024) relies on the $L_{2}$ norm of keys to determine token eviction. All these methods can be combined with Flash Attention. Notably, Devoto et al. (2024)’s approach demonstrates excellent granularity and controllability.
Merging, as a relatively recent method, represents a natural extension of Eviction, transitioning from hardmax to softmax. The current key challenges in this area include determining the merging region and selecting an appropriate scheme for combining keys and values.
Nawrot et al. (2024) proposed Dynamic Memory Compression (DMC), which use a single, after-trained dimension in key to determine whether to merge or not, and use attention score as the reference of merging. Wang et al. (2024) proposed the KVMerger, to merge KV by Gaussian weights and attention score.
Yu et al. (2024) proposed a novel approach to merge KV Cache of each token across different attention head, with a small amount of additional training, eventually turning an MHA model into a GQA model. Also requires additional training, Pang et al. (2024) proposed Anchor-LLM, which utilizes an innovative anchor-based self-attention, to help model learn how to merge KV Cache into a few token when training, achieving an outstanding performance.
@@ 5.2 Quantization
Another commonly utilized method for compressing Key-Value (KV) cache is through quantization. This approach effectively compresses data by mapping tensor values, originally in full precision, to discrete levels and storing them at a reduced precision. There are two primary categories of KV Cache quantization: Full quantization and KV Cache-only quantization. Full quantization involves compressing both the model weights and the KV Cache, thereby reducing the memory footprint of the entire model. On the other hand, KV Cache-only quantization specifically targets the KV Cache activations, selectively compressing them to conserve memory while leaving the model weights in their original state. This tailored approach can offer a balance between model efficiency and performance.
Figure 3: The comparison between different KV Cache policies
KV Cache only quantization
@@
Dong et al. (2024a) introduces LESS (Low-rank Embedding Sidekick with Sparse policy), which integrates a constant-sized cache with eviction-based cache methods (Section 5). Inspired by recurrent networks, LESS utilizes a constant-sized low-rank cache by replacing the softmax with a separable similarity metric calculated by learnable row-wise functions. By doing so, LESS obtains the ability to accumulate history token information before they are discarded from KV Cache, allowing for continued access to previous information.
Yang et al. (2024b) proposd Mixed-precision KV Cache (MiKV), a reliable cache compression method that retains evicted KV pairs in reduced precision to preserve information and maintains important KV pairs in higher precision to ensure generation quality. MiKV employed Dynamic Outlier Awareness to dynamically balance the outliers manifested in the query and keys to reduce quantization error. Dynamic Outlier Awareness multiplies and divides a channel balancer to the keys and queries to mitigate the impact of outliers.
@@
Yue et al. (2024) took FlexGen a step further and proposed WKVQuant. For weight quantization, WKVQuant ultize the OmniQuant (Shao et al., 2023). For KV Cache quantization, WKVQuant employs following three strategies: Past Only Quantization, Two-dimensional Quantization and Cross-block Reconstruction Regularization. Different from KVQuant (Hooper et al., 2024) which focus on calibrating for per-channel(token) thresholds to ignore outlier, WKVQuant focuses on aligning and smoothing each channel and token.
@@ 6 Evaluation
In this section, we first introduce the datasets commonly used in KV Cache optimization, followed by a presentation of frequently used evaluation metrics.
@@ 6.1 Datasets
Long context benches
Currently, there are several benchmarks that rely on extracting information from extremely long texts to answer questions. These benchmarks typically require models to read a substantial reference document, often containing 4,000 to 20,000 tokens, and then respond to questions. In some cases, the longest texts may even approach 200,000 tokens. Typical examples are LongBench (Bai et al., 2023), ZeroSCROLLS (Shaham et al., 2023), L-Eval (An et al., 2023), BAMBOO (Dong et al., 2024c), XL2bench(Ni et al., 2024), InfiniteBench (Zhang et al., 2024b), and LooGLE (Li et al., 2023). However, these benchmarks face challenges related to model-specific knowledge requirements, fixed text lengths that are difficult to adjust, and the uncontrollable and subjective nature of content generated by the models.
Key retrieval
The other method for evaluating the ability of models to handle extremely long texts involves directly extracting a specified key from unrelated context. This approach is simpler, more flexible, and more controllable. Importantly, it does not rely on the model’s existing knowledge. Examples are Needle in a Haystack (Kuratov et al., 2024), Passkey Retrieval (Mohtashami & Jaggi, 2023), and RULER (Hsieh et al., 2024). However, critics argue that this method is overly simplistic. Examples of these tasks can be found in Appendix B Figure 4 and Figure  5.
@@
In addition to datasets specifically designed for long texts, the length of traditional shorter test sets can be extended through a few-shot (Brown et al., 2020) format or by simulating multi-turn dialogues, in order to test the model’s capabilities with long texts. However, in long-text tests based on few-shot, texts at a greater distance only serve as a reference and do not contain information that the model needs to extract or conditions for inference, so the reference value of the results is limited.
@@ 6.2 Evaluation Metric
For KV Cache, the most intuitive optimization indicator would be the memory space occupied by each token. The LLaMA2-7B model, as a typical example, theoretically occupies $0.5$MB of memory for each KV Cache entry. Please note that when measuring this indicator rigorously, the fragment space generated by the token should also be taken into account, that is, it is best to measure according to the actual memory occupancy, rather than a value calculated through a string of parameters.
@@
Throughput and latency are important indicators for measuring the time efficiency of a model. Throughput, usually measured in tokens per second (token/s), represents how many new tokens the model can generate per second. The higher the throughput, the better the model efficiency. In the Decoding phase, latency is usually considered to be the time required to generate each new token, which is the reciprocal of the throughput, typically in milliseconds.
@@ 7 Key Takeaways
This review, following the footsteps of prior work, takes a closer look into KV Cache optimization, uncovering its complexities and proposing strategies for its optimization. We hope to propose some insights and suggested directions for future research that are not just reflections of current trends but also an invitation to explore the uncharted territory of LLMs.
Principles of KV Cache Optimization: At the core of optimizing KV Cache lies the principle of reducing memory consumption. This can be achieved by further compressing ’K’ (Keys) or ’V’ (Values) in the KV pairs. Techniques to compress these components directly impact the efficiency of the models, especially in terms of memory usage and processing speed.
Trade-offs in Deletion vs. Compression: Whether to delete less important KV pairs to save memory or to compress the KV Cache without deletion remains an open question. While deletion might offer immediate memory relief, it could potentially compromise the model’s performance. In contrast, better compression techniques strive to retain information integrity while reducing memory footprint.
Extremes in KV Cache Management: A more radical approach could involve storing the KV Cache externally, possibly on a different storage medium. This method would transform KV Cache management into a retrieval challenge, where the relevant KV pairs are fetched and reintegrated into the model as needed. While this could reduce memory usage on primary devices, it would introduce complexities in retrieval and integration processes.
Future Directions in Storage and Retrieval Technologies: These discussions point towards an evolving future where storage and retrieval technologies might become as crucial as the computational models themselves. Innovations in how KV Cache is managed, stored, and accessed could open up new avenues for making LLMs more efficient and versatile.
@@ 8 Conclusion
In this review, we highlighted the significance and the multifaceted nature of optimizing KV Cache in Large Language Models (LLMs). This review traversed through a variety of methods, from the training and deployment stages to post-training optimizations. Each method, whether it be the architectural changes in the pre-training phase, framework optimizations during deployment, or dynamic strategies like eviction and quantization in the post-training phase, offers unique insights into mitigating the challenges posed by the KV Cache’s memory-intensive nature. Moreover, the exploration of different evaluation metrics for long-text performance underpins the importance of a balanced approach that considers both efficiency and model capabilities. As the field continues to evolve, it is clear that the optimization of KV Cache will remain a critical area of focus, offering a pathway towards more efficient and environmentally responsible use of LLMs. We hope this review can be a roadmap to guide learners venturing through this dynamic and rapidly progressing field.
@@ Appendix A Popular Models With GQA
  • $S_{HK}$ and $S_{KV}$: Embedding size (in bytes, bfloat16 as data type) and corresponding KV Cache increment per token.
@@
  • $\mathcal{R}$: Represents the ratio of the KV Cache size generated for each token to the size of the embedding vector. A higher ratio indicates that the information for this token has been expanded by a greater factor. Notably, when using the GQA method, this value is significantly lower compared to non-GQA methods, implying that the information produced is relatively compressed.
@@ Appendix C Other Methods
In addition to the above methods, there are also some methods that modify the Attention mechanism to completely eliminate KV Cache. Furthermore, by directly compressing the prompt or the embedding, the length of KV Cache can also be reduced by decreasing the number of tokens.
@@ C.1 Linear-Transformers
Katharopoulos et al. (2020); Wang et al. (2020); Verma (2021) proposed a linear attention mechanism. By removing the softmax function and slightly adjusting $W_{\{q,k,v\}}$, the computational complexity of the Attention part in the Transformer is transformed from quadratic to linear, thereby achieving an inference speed consistent with Recurrent Neural Networks (RNNs), and totally eliminate the KV Cache. Other popular Linear-Transformers, mainly RNN-like or SSM models, includes RetNet (Sun et al., 2023b), RWKV (Peng et al., 2023b), Mamba (Gu & Dao, 2024), etc.
Zhang et al. (2024a); Arora et al. (2024) are some improvements to linear attention. These include the use of naive Attention mechanisms at close range, and the optimization of the performance of linear Transformers through Spiky and Monotonic Weights.
@@ C.2 Prompt and Embedding Engineering
LLMLingua (Jiang et al., 2023a) uses a budget controller, an iterative token-level compression algorithm and a small model to estimate which token is important. LongLLMLingua (Jiang et al., 2023b) further improves LLMLingua in long context scenarios.